OpenStudy (anonymous):

simplify the expression (root-1)/(2-4i)-(5+3i)

OpenStudy (anonymous):

answer choices are:

OpenStudy (anonymous):

\[\frac{ i }{ (2-4i)-(5+3i) }\]?

OpenStudy (anonymous):

@Hero

hero (hero):

@doulikepiecauseidont has it right. Simplify the denominator now.

OpenStudy (anonymous):

how do i simplify it if there are negative numbers on the bottom?

hero (hero):

(2 - 4i) - (5 + 3i) Remember `-(5 + 3i)` = `-1(5 + 3i)` = -5 - 3i

OpenStudy (anonymous):

how can that be the same thing if one sign is the same and the other is opposite?

hero (hero):

Then you will have 2 - 4i - 5 - 3i When you re-arrange it by placing like terms together you have 2 - 5 - 4i - 3i Afterwards, you can re-group things again: (2 - 5) - (4 + 3)i or - ( 5 - 2) - (4 + 3)i = -3 -7i

OpenStudy (anonymous):

is -3-7i the same as -7-3i?

hero (hero):

NO

OpenStudy (anonymous):

is it the same as 3+7i?

hero (hero):

no

hero (hero):

Basically, this just means you have no clue what to do next to simplify this: \[\frac{i}{-3 - 7i}\]

OpenStudy (anonymous):

this is a valid statement

hero (hero):

You can simplify it further to get \[-\frac{i}{3 + 7i}\]

hero (hero):

But you have to multiply top and bottom by the conjugate in order to continue simplifying.

OpenStudy (anonymous):

so I would then multiply the top and bottom by (3-7i)

hero (hero):

Yes. Make sure you do so carefully.

OpenStudy (anonymous):

so the top would be 3i-7i^2 and the bottom would be 9+49i^2 ?

hero (hero):

The bottom will be a difference of squares. But you don't do it like that.

OpenStudy (anonymous):

so I don't foil and cancel the middle?

hero (hero):

\[-\frac{i(3-7i)}{(3 + 7i)(3 - 7i)}\] If you did it correctly, you'd get one of your answer choices.

OpenStudy (anonymous):

\[i^2=-1\]

OpenStudy (anonymous):

okay the top is 3i*-7i^2 which simplifies to 3i+7, yes?

hero (hero):

Yes, but don't forget about the negative in front of the fraction

hero (hero):

You have to distribute it across the numerator expression

OpenStudy (anonymous):

which means that the numerator is really -3i+7?

hero (hero):

No, not both

hero (hero):

You don't know how to distribute a negative? -(3i + 7) = ?

OpenStudy (anonymous):

sorry, typo, meant to say -3i-7

hero (hero):

What about the bottom? You should have gotten: \(9 - 49i^2\).

hero (hero):

And after replacing \(i^2\) with -1, you should have \(9 - 49(-1))\)

OpenStudy (anonymous):

ooooooh, revelation. I got it now, thanks soo much, both of you!

hero (hero):

Just to confirm, what is the final result?

OpenStudy (anonymous):

(-7-3i)/58

OpenStudy (anonymous):

@Hero do you know anytihng about bearings?

hero (hero):

I know about bearings

OpenStudy (anonymous):

can you help me with my question, it says the play is travelling on a bearing of 340 degrees at 325 mph, doesn't that mean they are in QII not QIV?

OpenStudy (anonymous):

the plane*

OpenStudy (anonymous):

@Hero

hero (hero):

Negative. You need to brush up on your quadrants. The plane is traveling in the 4th Quadrant.

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

 isn't that a bearing of 340 degrees(20 degrees from the north line)